OpenStudy (anonymous):

a car travels at an average speed of 52 miles per hour,how many miles does it travel in 3 hours and 30 minutes?

OpenStudy (anonymous):

how do i solve this

OpenStudy (amistre64):

well, how many miles does it travel in 1 hour?

OpenStudy (anonymous):

i dont know?

OpenStudy (amistre64):

"a car travels ... 52 miles per hour"

OpenStudy (anonymous):

ok ...

OpenStudy (amistre64):

so, how many miles does it go in 1 hour?

OpenStudy (anonymous):

52?

OpenStudy (amistre64):

good and if it goes 52 miles in one hour, then after 3 hours, how far has it traveled? 52+52+52 = ??

OpenStudy (anonymous):

156

OpenStudy (amistre64):

and since 30 minutes is half an hour, its safe to assume that it only went half of 52 miles for the remaining part 156 + (half of 52)

OpenStudy (anonymous):

26

OpenStudy (amistre64):

yes, 156 + 26 = 182 to use a formula: distance = "speed" times "time" distance = 52 * 3.5 = 182

OpenStudy (amistre64):

.... youre welcome? and good luck :/